# PointCompress3D - A Point Cloud Compression Framework for Roadside LiDARs in Intelligent Transportation Systems

### v1

@@ IV-A1 PSNR
IV-A2 Chamfer Distance
It is similar to PSNR in terms of a measurement system for reconstruction error. The Chamfer distance between two point clouds $P_{1}=\left\{\bm{x_{i}}\in\mathbb{R}^{3}\right\}_{i=1}^{n}$ and $P_{2}=\left\{\bm{x_{j}}\in\mathbb{R}^{3}\right\}_{j=1}^{m}$ is defined as the average distance between pairs of nearest neighbors between $P_{1}$ and $P_{2}$ i.e.
where $\mathrm{NN}(\bm{x},P)=\operatorname{argmin}_{\bm{x^{\prime}}\in P}\left\|\bm{x}-\bm{x^{\prime}}\right\|$ is the nearest neighbor function and $n,m$ are the number of points in $P_{1}$, $P_{2}$ respectively. The Chamfer distance value has a minimum of 0, meaning that the two sets are spatially identical. We can also adjust the distance metric |||||$\cdot$||| | ⋅ | | with various distance metrics such as Manhattan or Euclidian distance metric.
IV-A3 BPP (bits per point)
The compression rate is measured using bits per point (BPP), indicating the bits needed to store a single data point. For instance, a binary pcd file with 16-bit precision for $(x,y,z)$ coordinates requires $3\times 16=48$ bits, which equals 6 bytes. In contrast, ASCII format uses 3 to 36 bytes due to its readable text storage (1 byte per character). Compression formats achieve significantly smaller sizes (refer to Table I).
IV-A4 Metric Graph
As we can see in Table I, for each bits per point value representing compression method efficiency, there will be a PSNR value indicating the reconstruction quality of that specific BPP from a certain compression method. We can control BPP indirectly in several ways, including quantization level, octree depth, or entropy coding efficiency. This will also indirectly affect the PSNR value. It is important to note that the BPP value is proportional to PSNR. When the BPP value is higher, PSNR will also be higher and vice versa. Therefore, we usually use a 2D graph to map the efficiency and effectiveness of compression methods and compare their performance using those graphs.
IV-A5 Compression Speed
We also measure the speed of compression and decompression individually to run each operation. The speed is calculated only for compression and decompression processes, neglecting the speed required to write or read data from storage.
IV-A6 Memory Consumption
Finally, we report the allocated memory on the GPU (VRAM) for each compression algorithm. We measure the memory consumption separately for the encoding and decoding steps.
IV-A7 Mean Average Precision
The mean average precision (mAP) metric quantifies the average precision scores across multiple object categories, evaluating a detection model’s performance across different classes. For our original uncompressed point cloud data we get an $mAP$ of 20.11
IV-B Experiment Setup
Datasets We use the TUMTraf Intersection dataset [27] (release R02, sequence S02) and the TUMTraf V2X Cooperative Perception [29] dataset. Both were labeled with the 3D bounding box annotation tool (3D BAT) [39, 29]. The latter one contains ten 10 s long sequences. Each sequence has 100 LiDAR point cloud scans and their corresponding 2D projection image (range, reflectivity, signal, near-infrared).
Hardware Setup Our framework runs inside a docker container with Docker version 20.10.20 under Ubuntu 20.04. We use 3x NVIDIA GeForce RTX 3090 GPUs, each having 24576 MiB of VRAM.
IV-C Quantitative Results
In this section, we evaluate different methods on our TUMTraf datasets and provide quantitative results for bits per point, PSNR d1 and d2, and codec [^3] The codec speed is given for the encoding and decoding process separately. speed.
TABLE III: Point cloud image compression size (in KB).
@@ IV-C Quantitative Results
TABLE IV: Evaluation of the compression and the detection performance based on different input representations (3D to 2D transformation).
@@
Figure 6: We show the ablation study results for Depoco of reconstructed point clouds for different point and grid sizes on the TUMTraf V2X Coop. Perception [29] dataset.
@@
Figure 7: Qualitative reconstruction results of Depoco, 3DPCC and Draco on the TUMTraf V2X Coop. Perception [29] dataset.
Referring to Sec. III-B1, we find that the max. number of points and grid size are the most impactful parameters for Depoco. Tuning them further, we present the results in Table II. The decoding speed remains consistent across configurations, while the encoding speed, GPU usage, and BPP increase with higher max. number of points. We calculate the BPP by dividing the encoded binary file size by the number of points in the original point cloud. We proceed by analyzing the conversion of 3D point clouds to 2D images. Each 3D point cloud frame yields three images (AIR): azimuth, intensity, and range. An example result is shown in Fig. 9. The transformation speed from 3D to 2D points is 300 ms, and from 2D to 3D points is 19 ms, while the ImageMagick PNG to JPG conversion takes approximately 15 ms. The average size of the range, azimuth, and intensity images from conversion compared to the original ones can be seen in Sec. IV-C. Further, we compress the images using PNG compression using the PIL python library, online compression TinyJPG [40], and ImageMagick library with compression level 100 for the range image only. As we can see from the table, the PNG compression by PIL hardly compresses the images followed by the ImageMagick, while the TinyJPG performs really well. At last, we evaluate the Draco method, achieving encoding and decoding times of 60 ms and 80 ms, a PSNR of 52 dB and 144 dB, and an mAP of 20.01 and 20.11, respectively. Finally, we compare Depoco’s performance to 3DPCC and Draco, visualizing it in a 2D graph (see Fig. 8).
@@
Figure 8: From left to right: Visualization of BPP and PSNR metrics for Depoco, 3DPCC and Draco.
@@
Figure 9: Result of the 3D to 2D transformation. From top to bottom: range, azimuth, and intensity image. We post-process the images and show the front part of the LiDAR.
IV-D Qualitative Results
@@
Here, we show the reconstruction result for different parameters of Depoco as shown in Tab. II. Furthermore, we provide qualitative results for the reconstruction from range images as depicted in Tab. IV. We illustrate how the PSNR value corresponds to the qualitative result. As shown in Fig. 6, increasing the max. number of points parameter will add more points to the scene after decoding. Note that we keep the grid size low, either 8x8x3 or 16x16x6, to get a defined object. Otherwise, the object will be blurry, as we can see in the 200k|24x24x9 setting, and it will get even more blurry as the grid size gets larger. Keeping the grid size small also keeps the PSNR higher, meaning that the point cloud reconstructed looks more similar to the original one. Fig. 7 illustrates the optimal reconstruction quality achieved at minimal bits per pixel for Depoco, 3DPCC, and Draco.
V Conclusion & Future Works
@@
After evaluating the three methods, Draco emerges as the most promising method, with encoding and decoding speeds exceeding 10 FPS alongside good PSNR and mAP values. Draco’s PSNR and mAP values consistently correlate, with higher PSNR values leading to better mAP values. Quantization bits are revealed as the most influential parameters in Draco, with value of 16 yielding mAP values of 20.01 and 20.11 when set to the maximum of 30 bits. On the other hand, Depoco faces speed issues despite configurations surpassing the baseline mAP of 20.11. The surpassing result is due to Depoco’s introduction of additional points, enhancing object recognition but leading to inconsistent PSNR-mAP relationships. Lastly, 3DPCC exhibits a good PSNR value without compression (36.23 vs. 32.88 for Depoco) but a very poor mAP of only 7.49, leaving no option for improvement except fixing the transformation to prevent loss of point clouds as what we have found. Future work includes deploying our framework on the live system, and the extension to other LiDAR types and manufacturers.

### v2

@@ 4 Evaluation
Table 2: Parameter tuning results for Depoco on the max. number of points and the grid size. We use the TUMTraf Intersection dataset to find the best parameters and evaluate the compression method on the TUMTraf V2X Cooperative Perception dataset.
4.1 Performance Metrics
In compression, we encounter diverse metrics assessing performance and quality. We balance speed, quality, and performance and determine the best trade-off. We assess the reconstruction quality at specific compression rates, comparing metric graphs to determine method superiority.
@@ 4.1 Performance Metrics
Table 3: Results of the point cloud image (range, azimuth, and intensity) compression of 3DPCC on the TUMTraf-I dataset. Size is given in KB.
PSNR. Following [19] based on [34], we use PSNR (Peak Signal-to-Noise Ratio) with the d1 and d2 metric, as reconstruction metric. PSNR is a widely used metric in image and video compression to evaluate the quality of the compressed image or video compared to the original ones. It measures the reconstruction quality in both directions by comparing the mean squared error (MSE) between the original and compressed point cloud (see Eq. 1). The PSNRs of the two directions are then combined to obtain a single symmetric PSNR value with the maximum pooling function, as defined in Eq. 2.
$\text{PSNR}_{A,B}=10\cdot\log_{10}\left(\frac{{\text{p}^{2}_{s}}}{{d^{MSE}_{A,B}}}\right)$ (1)
$\text{PSNR}=max(\text{PSNR}_{A,B},\text{PSNR}_{B,A})$ (2)
where $p$ is the signal peak $d$ is the average squared error (i.e. MSE) between all point in point cloud A and their corresponding neighbor in point cloud B. PSNR is measured in decibels (dB), and a higher PSNR value indicates better quality because the MSE is smaller. In the case of point clouds, the MAX value can be set in various ways, depending on the compression type, such as voxelized or not [19]. The most common method is to set it to the value of the maximum diagonal of the box bounding of the point cloud.
@@
Table 4: Evaluation of the 3DPCC compression and the PointPillars 3D object detection performance on the TUMTraf-I dataset. We compare four different input representations.
@@
Figure 5: Visualization of the ablation study results for Depoco. We show the reconstructed point clouds for different grid sizes (x-axis) and point sizes (y-axis) on the TUMTraf V2X Cooperative Perception [49] dataset and report the PSNR d1 and d2 metrics.
@@
Figure 6: Qualitative reconstruction results. Qualitative comparison of Depoco, 3DPCC and Draco on the TUMTraf V2X Cooperative Perception [49] dataset. Draco outperforms Depoco and 3DPCC regarding the BPP and PSNR d2 metrics.
Chamfer Distance. It’s similar to PSNR in terms of a measurement system for reconstruction error. The Chamfer distance between two point clouds $P_{1}=\left\{\bm{x_{i}}\in\mathbb{R}^{3}\right\}_{i=1}^{n}$ and $P_{2}=\left\{\bm{x_{j}}\in\mathbb{R}^{3}\right\}_{j=1}^{m}$ is defined as the average distance between pairs of nearest neighbors between $P_{1}$ and $P_{2}$ i.e.
where $\mathrm{NN}(\bm{x},P)=\operatorname{argmin}_{\bm{x^{\prime}}\in P}\left\|\bm{x}-\bm{x^{\prime}}\right\|$ is the nearest neighbor function and $n,m$ are the number of points in $P_{1}$, $P_{2}$ respectively. The Chamfer distance has a minimum of 0, meaning that two sets are spatially identical. We can also adjust the distance metric |||||$\cdot$||| | ⋅ | | with various distance metrics such as Manhattan or Euclidian distance metric.
@@
Figure 7: Visualization of BPP and PSNR metrics. From left to right: Depoco, 3DPCC and Draco. We calculate the Area Under the Curve (AUC) metric for different compression levels and see that Draco outperforms the other methods.
Bits per point. Compression rate is measured using bits per point (BPP), indicating the bits needed to store a single point. For instance, a binary pcd file with 16-bit precision for $(x,y,z)$ coordinates requires $3\times 16=48$ bits, which equals 6 bytes. In contrast, ASCII format uses 3-36 bytes due to its readable text storage (1 byte per character). Compression formats achieve significantly smaller sizes.
Metric Graph. For each bits per point value representing compression method efficiency, there will be a PSNR value indicating the reconstruction quality of that specific BPP from a certain compression method (see Tab. 1). We can control BPP indirectly in several ways, including quantization level, octree depth, or entropy coding efficiency. This will also indirectly affect the PSNR value. It is important to note that the BPP value is proportional to PSNR. When the BPP value is higher, PSNR will also be higher and vice versa. Therefore, we usually use a 2D graph to map the efficiency and effectiveness of compression methods and compare their performance using those graphs.
Compression Speed. We also measure the compression and decompression speed individually to evaluate the performance of the methods. The speed is calculated only for compression and decompression processes, neglecting the speed required for pre- or postprocessing steps.
Memory Consumption. Finally, we report the allocated memory for the encoding and decoding on the GPU (VRAM) for each compression algorithm.
Mean Average Precision. The mean average precision (mAP) metric quantifies the average precision scores across multiple object categories, evaluating a detection model’s performance across different classes.
4.2 Experiment Setup
Datasets. We use the TUMTraf Intersection [45] (release R02, sequence S02) and the TUMTraf V2X Cooperative Perception [49] datasets, labeled with 3D BAT [48, 49], for our experiments. The latter one contains ten 10 s long sequences. Each sequence has 100 LiDAR point cloud scans and their corresponding 2D projection images (range, reflectivity, signal, near-infrared).
Hardware Setup. Our framework runs inside a docker container under an Ubuntu operating system. We use 3x NVIDIA GeForce RTX 3090 GPUs for our experiments, each having 24576 MiB of VRAM.
4.3 Quantitative Results
In this section, we evaluate different methods on the TUMTraf datasets and provide quantitative results for bits per point, PSNR d1 and d2, encoding and decoding speed.
Figure 8: Qualitative reconstruction results of Depoco. Left to right: 1) Original and reconstructed point cloud of the TUM Traffic A9 Highway (TUMTraf-A9) dataset [8]. 2) Original and reconstructed point cloud of the TUM Traffic Intersection (TUMTraf-I) dataset [45].
We find that the max. number of points and grid size are the most impactful parameters for Depoco. We present the results of our parameter fine-tuning in Table 2. The decoding speed remains consistent across configurations, while the encoding speed, GPU memory usage, and BPP increase with a higher max. number of points. We calculate the BPP by dividing the encoded binary file size by the number of points in the original point cloud. We proceed by analyzing the conversion of 3D point clouds to 2D images. Each 3D point cloud frame yields three images (AIR): azimuth, intensity, and range. An example result is shown in Fig. 9. The transformation speed from 3D to 2D points is 300 ms, and from 2D to 3D points is 19 ms, while the ImageMagick PNG to JPG conversion takes approximately 15 ms. The average size of the range, azimuth, and intensity images from conversion compared to the original ones can be seen in Sec. 4.1. Further, we compress the images using PNG compression using the PIL python library, online compression TinyJPG [1], and ImageMagick library with compression level 100 for the range image only. As we can see from the table, the PNG compression by PIL hardly compresses the images followed by the ImageMagick, while the TinyJPG performs really well. At last, we evaluate the Draco method on the TUMTraf-I dataset, achieving encoding and decoding times of 60 ms and 80 ms, a PSNR of 52 dB and 144 dB, and an mAP of 20.01 and 20.11, respectively. Finally, we compare Depoco’s performance to 3DPCC and Draco, visualizing it in a 2D graph (see Fig. 7).
@@ 4.3 Quantitative Results
Figure 9: Results of the 3D to 2D transformation. From top to bottom: range, azimuth, and intensity image. We post-process the images and show the front part of the LiDAR.
4.4 Qualitative Results
@@
In Fig. 5 we show the reconstruction results for different parameters of Depoco that were fine-tuned and listed in Tab. 2. We illustrate how the PSNR value correlated to the grid size and the number of points. Increasing the max. number of points parameter will add more points to the scene after decoding and lead to a higher PSNR value. Note that we keep the grid size low, either 8x8x3 or 16x16x6, to get a good reconstruction result. Otherwise, reconstructed objects are blurry, as we can see in the 200k|24x24x9 setting, and it will get even more blurry as the grid size gets larger. Keeping the grid size small keeps the PSNR high, meaning that the reconstructed point cloud looks more similar to the original one. Fig. 6 illustrates the optimal reconstruction quality achieved at minimal bits per pixel for Depoco, 3DPCC, and Draco. Finally, we show qualitative reconstruction results of Depoco on the TUMTraf-A9 and the TUMTraf-I dataset (see Fig. 8).
5 Conclusion and Future Work
@@
Draco emerges as the most promising method, with encoding and decoding speeds exceeding 10 FPS alongside good PSNR and mAP values. Draco’s PSNR and mAP values consistently correlate, with higher PSNR values leading to better mAP values. Quantization bits are revealed as the most influential parameters in Draco, with values of 16 yielding mAP values of 20.01 and 20.11 when set to the maximum of 30 bits. On the other hand, Depoco faces speed issues despite configurations surpassing the baseline mAP of 20.11. The surpassing result is due to Depoco’s introduction of additional points, enhancing object recognition but leading to inconsistent PSNR-mAP relationships. Lastly, 3DPCC exhibits a good PSNR d2 value of 36.23 compared to 32.88 for Depoco on the TUMTraf-I dataset. On the other hand, reconstructed point clouds with 3DPCC lead to a very low 3D object detection performance (mAP of 7.49). We found that fixing the transformation to prevent loss of point clouds can improve that. Future work includes deploying our framework on the live system, and the extension to other LiDAR types and manufacturers.
@@ Acknowledgment
Supplementary Material
https://pointcompress3d.github.io/
Figure 10: Visualization of point cloud reconstruction results. We show further reconstruction results of Depoco on the TUMTraf-A9 and TUMTraf-I dataset. First row from left to right: 1) Original and reconstructed point cloud on the A9 Highway showing a truck, 2) Original and reconstructed point cloud at an intersection showing a pedestrian. Second row from left to right: 3) Original and reconstructed point cloud on the A9 Highway showing a van, 4) Original and reconstructed point cloud at an intersection showing a car.
Contents
[…]
Appendix A Task definition
Point cloud compression involves reducing the data size of point clouds, represented as sets of points $\{(x_{i},y_{i},z_{i})\}$ in space, while preserving their geometric and attribute information. The task aims to minimize the storage $S$ and transmission time $t$ requirements by finding a compact representation $C$ such that $|C|<<|P|$, where $P$ denotes the original point cloud data. The goal of efficient compression algorithms is to achieve high compression rates $r=\frac{|P|}{|C|}$ with minimal loss of information $\epsilon$, thus ensuring that the reconstructed point cloud $P^{\prime}$ closely approximates $P(d(P,P^{\prime})\leq\epsilon)$. Compression algorithms are applied in applications like autonomous driving, virtual reality, and geographic information systems, where large volumes of 3D data need to be processed and transmitted efficiently.
Appendix B Further related work
This section provides an overview of additional point cloud compression methods that have been proposed in the literature.
B.1 Traditional point cloud compression
Some of the most commonly used traditional point cloud compression algorithms include MPEG PCC [34] and AVS PCC [22]. MPEG PCC is a standardization effort by the Moving Picture Experts Group (MPEG) for point cloud compression. AVS PCC is a similar standardization effort by the Audio Video Coding Standard Workgroup of China (AVS). Traditional point cloud compression algorithms are based on geometry coding, attribute coding, and entropy coding techniques. Yu et al. [41] propose a regularized projection algorithm to construct a reliable prediction relationship in the predictive structure. A simplified geometry prediction technique is proposed based on the regularized projection pattern. [33] introduces two new encoding methods for point cloud compression, namely Alternate Depth Compression (ADC) and Log-Polar (LP). In [25] the authors propose SCP, a model-agnostic spherical coordinate-based point cloud compression method. Moreover, a multi-level Octree is introduced to mitigate the reconstruction error for distant areas in the point cloud.
B.2 Learning based compression
Learning-based point cloud compression algorithms have shown promising results in terms of compression efficiency and reconstruction quality. EHEM [35] proposes a hierarchical attention structure that has a linear complexity to the context scale and maintains the global receptive field. Furthermore, a grouped context structure is presented to address the serial decoding issue caused by the auto-regression mechanism. PIVOT-Net [28] introduces a heterogeneous point cloud compression framework that unifies typical point cloud representations such as point-based, voxel-based, and tree-based representations. This allows compressing point clouds at different bit-depth levels. SparsePCGC [38] is a sparse point cloud geometry compression method that only performs convolutions on sparsely distributed most-probable positively occupied voxels.
Appendix C Implementation details
This section provides additional implementation details of our point cloud compression framework PointCompress3D.
C.1 Hardware and software setup
We run our compression framework on a GPU cluster with an AMD EPYC 7001 series CPU, 128 GB of RAM, and 3 x NVIDIA RTX 3090 GPUs. The framework runs in a Docker container based on the Ubuntu 20.04 operating system and is implemented in Python.
C.2 Parameter tuning
We perform parameter tuning to optimize the performance of our point cloud compression framework. The parameters include the grid size, the max. number of points, min. kernel size, kernel radius (encoding/decoding), quantization bits and compression level. We use grid search to find the optimal parameters for each model. The parameters are tuned on the TUMTraf-I dataset and are then used for testing the models on the TUMTraf-V2X dataset.
Appendix D Point Cloud Compression Datasets
This section provides an overview of the most commonly used point cloud datasets for evaluating point cloud compression algorithms [22].
  1. 1.
    KITTI [13]: Contains point clouds captured by a Velodyne LiDAR sensor mounted on a moving vehicle.
  2. 2.
    ShapeNet [7]: Consists of 3D models from various categories, including cars, chairs, and tables.
  3. 3.
    ModelNet40 [40]: Contains 3D models from 40 categories, such as airplanes, chairs, and tables.
  4. 4.
    SemanticKITTI [5]: A large-scale dataset of point clouds with semantic annotations.
  5. 5.
    TUMTraf-I [45]: Contains point clouds captured by two Ouster LiDAR sensors mounted on roadside infrastructure.
Appendix E Metrics
The metrics for evaluating point cloud compression performance and reconstruction quality can be categorized into performance metrics and quality metrics [33].
E.1 Performance metrics
  1. 1.
    Bit per point (BPP): Measures the size in bits used to represent a single point.
  2. 2.
    Bit-rate (BR): Total size of the computed frames divided by the frame transmission rate.
  3. 3.
    Compression ratio (CR): Ratio between the BPP of the raw point cloud and the BPP of the encoded point cloud.
  4. 4.
    Encoding and decoding time (ET, DT): Time taken to process the data and generate the corresponding output.
E.2 Quality metrics
  1. 1.
    Peak Signal-to-Noise Ratio (PSNR):
    • PSNR-D1: Intra-point MSE error calculated between the original and reconstructed point cloud.
    • PSNR-D2: Similar to PSNR-D1 but measures distances with respect to a plane.
  2. 2.
    PointSSIM (PSSIM): Measures perceptual degradations of point cloud data using statistical dispersion measurements for attributes such as colors and geometry.
  3. 3.
    Video Quality Assessment Point Cloud (VQA-PC): Uses spatial and temporal features extracted from a point cloud to estimate quality levels through a machine learning model.
  4. 4.
    Multi-Modal Point Cloud Quality Assessment (MM-PCQA): The point cloud is split into various 3D sub-models and rendered into 2D image projections. These projections are encoded with neural networks, and the quality level is estimated through a quality regression block.
These metrics are used to evaluate both the efficiency of the compression process and the quality of the reconstructed point cloud.
Appendix F Detailed visualization results
We provide additional qualitative results of the point cloud reconstruction process using Depoco on the TUMTraf-A9 and TUMTraf-I datasets. Figure 10 shows the original and the reconstructed point cloud on the A9 Highway with a truck and a van. Moreover, we show the original and reconstructed point cloud at an intersection showing a pedestrian an a car. These results demonstrate the effectiveness of Depoco in reconstructing point clouds with different objects and scenes in the TUMTraf datasets.
Appendix G Failure cases and limitations
G.1 Failure cases
Learning-based point cloud compression algorithms like 3DPCC [4] can encounter several failure cases. These include overfitting to the training data, resulting in poor performance on unseen data, and underfitting, where the model fails to capture underlying patterns. They may also struggle with data imbalance, performing poorly when one class is overrepresented, and with outliers, which can negatively impact the model’s performance on datasets that contain atypical data points.
G.2 Limitations
Depoco [39] is no longer state-of-the-art (SOTA) and has only been evaluated on the SemanticKITTI dataset. It is outperformed by Draco on the TUMTraf-I dataset and does not consider the compression of point attributes. Similarly, 3DPCC [4] also does not consider the compression of attributes except the intensity and XYZ coordinates. It requires calibration data, range images, and azimuth maps for training, which are not always available in practice. It is also outperformed by Draco [12] on TUMTraf-I and is prone to overfitting. Finally, Draco is not as efficient as 3DPCC or Depoco in terms of decoding point clouds.